Northwood Company manufactures basketballs. The company has a ball that sells for $25. At present, the ball is manufactured in a small plant that relies heavily on direct labor workers. Thus, variable expenses are high, totaling $15.00 per ball, of which 60% is direct labor cost.

Last year, the company sold 60,000 of these balls, with the following results:

Sales (60,000 balls) $ 1,500,000
Variable expenses 900,000
Contribution margin 600,000
Fixed expenses 375,000
Net operating income $ 225,000

Required:

1. Compute (a) last year's CM ratio and the break-even point in balls, and (b) the degree of operating leverage at last year’s sales level.

2. Due to an increase in labor rates, the company estimates that next year's variable expenses will increase by $3.00 per ball. If this change takes place and the selling price per ball remains constant at $25.00, what will be next year's CM ratio and the break-even point in balls?

3. Refer to the data in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next year to earn the same net operating income, $225,000, as last year?

4. Refer again to the data in (2) above. The president feels that the company must raise the selling price of its basketballs. If Northwood Company wants to maintain the same CM ratio as last year (as computed in requirement 1a), what selling price per ball must it charge next year to cover the increased labor costs?

5. Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40.00%, but it would cause fixed expenses per year to double. If the new plant is built, what would be the company’s new CM ratio and new break-even point in balls?

6. Refer to the data in (5) above.

a. If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $225,000, as last year?

b. Assume the new plant is built and that next year the company manufactures and sells 60,000 balls (the same number as sold last year). Prepare a contribution format income statement and Compute the degree of operating leverage.
Answer #1

CM ratio = CM/Sales

= 600,000/1,500,000

=40%

Break point in Balls = Fixed costs/CM per unit

= 375,000/10

= 37,500 units

Degree of operating leverage = CM/Operating income

= 600,000/225,000

=2.6667

2.CM Ratio = (25-15-3)/25 = 28%

Break even point = 375,000/7 = 53,571.43 balls

3.Desired operating income = $225000

Add: Fixed costs = 375000

Desired contribution margin = $600,000

Number of balls = 600,000/7 = 85,714.29 balls

4.Selling price = Variable cost/Variable cost ratio

= 18/0.6 = $30

5.CM Ratio = (25-15*0.6)/25 = 64%

Break even point = 375,000*2/16

= 46,875 balls

6a. Balls to be sold = (225,000+750,000)/16

= 60,937.5 balls

b.Income Statement

Sales 60,000*25

1,500,000

Variable expenses 60,000*9

540,000

Contribution Margin

960,000

Fixed costs

750,000

Operating income

210,000

DOL = 960,000/210,000

= 4.5714